Second Beamer Example

Till Tantau May 5, 2006

1 The first section

This is the first section of the article version. In the presentation, there is a frame containing an overlay. The exact two slides of this overlay are shown in Figures 1 and 2.

.epsi .eps .pspgflastimage

Figure 1: The first slide. Note the partly covered second item.

.epsi .eps .pspgflastimage

Figure 2: The second slide. Now the second item is also shown.

We can also include the frame in the article version “just like this”: This is a frame with two overlays.

• The first item…

• … and the second one.

We could have suppressed the frame in the article version by adding the overlay specification

<presentation>.